One sample has a mean of M = 6 and a second sample has a mean of M = 12. The two samples are combined into a single set of scores.

(b) What is the mean for the combined set if the first sample has n = 4 scores and the second sample has n = 8?


(c) What is the mean for the combined set if the first sample has n = 8 scores and the second sample has n = 4?
Answer #1

Here we have :

A) The mean will be equal to
X = (6*4+12*8)/(4+8) = 10
B) The mean of the combined set would be
X = (6*8 + 12*4) / (4+8) = 8